Tom is 6 times as old as Todd. In 3 years Tom's age will be 33 years more than 2 times as old as Todd. How old is Tom?
Answer by mananth(16946)

Tom is 6 times as old as Todd. In 3 years Tom's age will be 33 years more than 2 times as old as Todd. How old is Tom?
...
Todd be x years
Tom = 6x
..
after 3 years
Todd's age = x+3 years
Tom's age = 6x+3 years
..
6x+3 = 33+2(x+3)
6x+3=33+2x+6
6x+3=2x+39
-3 -3
6x=2x+36
-2x -2x
6x-2x=36
4x=36
/4
x= 9 years..... Todd's age
Tom's age = 6x = 6*9 = 54 years
...
m.ananth@hotmail.ca